# Bifurcations and canards in two coupled FitzHugh–Nagumo equations July 25, 2025

### v1

@@ 5.2. Double fold points
For (b), we have already established in Subsection 5.1 that in this case $\det DH\left(x_{1}^{*},x_{2}^{*}\right)>0$. From the expression (12) and assertion (a) it follows that when $c$ tends to $2\sigma/\sqrt{3}$ then $\det DH\left(x_{1}^{*},x_{2}^{*}\right)$ tends to 0. Therefore, for $c$ close to $2\sigma/\sqrt{3}$ we have $0<\det DH\left(x_{1}^{*},x_{2}^{*}\right)<1/4$ and since $\operatorname{Tr}DH\left(x_{1}^{*},x_{2}^{*}\right)=1$, the point $\left(x_{1}^{*},x_{2}^{*}\right)$ is a node. ∎
@@ 6. Discussion
This article is an analysis of the dynamics of two identical FitzHugh–Nagumo equations symmetrically coupled through the slow equations. In this section we compare our results to findings by other authors using different types of coupling.
@@
We obtain persistent synchronous (i.e. symmetric) solutions for an open set of parameters. We also show that approximate synchrony will persist if the symmetry is broken either by taking slightly different coupling constants or by a small change in the equations governing one of the cells. Synchronised periodic solutions were also observed analytically by Campbell and Waite [2] and numerically in passing by Hoff et al. [14] when two FHN were coupled symmetrically and bidirectionally through the fast equations, but the latter do not report on persistence under symmetry breaking, their focus being on chaotic solutions. Kawato et al. [15] obtain both the synchrony and its persistence for two FHN with simultaneous slow-to-slow and fast-to-fast connection. Pedersen et al. [22] provide bifurcation diagrams in the synchrony plane for two FHN coupled symmetrically and bidirectionally through the fast equations, showing steady-state and Hopf bifurcations that create both stable and unstable limit cycles.
@@
The FitzHugh–Nagumo equations are symmetric when $c=0$, this provides an additional symmetry to the coupled equations, with a flow-invariant fixed point subspace. From this we obtain the coexistence of different stable solutions and we show that it persists when the additional symmetry is broken for small $c\neq 0$. The additional symmetry does not hold when the equations are coupled through the fast equations, so this bistability is a characteristic feature of the type of coupling analysed here.
@@
Bistability has been found in the work by Kawato et al. [15] coupling together the slow variables and the fast variables to each other in two FHN systems. They have obtained the coexistence of synchrony and antisynchrony through Hopf bifurcation. It has also been found by Campbell and Waite [2] who considered electrical coupling of two FHN models through the fast equation although none of these authors used the fast-slow structure. In [2] it is shown that when the magnitude of the coupling is strengthened, periodic orbits undergo several bifurcations (namely resonant Hopf-Hopf interactions) leading to the coexistence of a chaotic attractor with an attracting periodic orbit.
@@
The symmetry in the model we have analysed limits the possible dynamic outcomes, and yet it still allows a number of interesting features. The presence of a canard induces small amplitude symmetry breaking oscillations before a solution converges to a synchronised periodic solution of large amplitude, shown in Figure 6. Santana et al. [24] report “transient chaos”, a complicated and long transient, in their numerical description of two FHN with different parameter values coupled through the fast equations. Canard-induced small amplitude transients were also found by Krisitiansen & Pedersen [16] in two identical FHN coupled through the fast equations. In their case the attracting periodic solution does not lie in the synchrony plane. The small amplitude transients arise from a canard at a cusp point in the critical manifold, in our case this manifold does not have cusps, so the origin of the transient oscillations is not the same, here they arise at a folded node.
Sustained mixed-mode oscillations arising from canards, like those in Figures 7 and 8, are ubiquitous in coupled FHN. They appear in asymmetric coupling of the fast equation to the slow one in Doss-Bachelet et al. [5] and in Krupa et al. [18] where three different time scales are considered. Desroches et al. [4] find them in self-coupled FHN analogous to unidirectional coupling through the fast equation. Krupa et al. [17] find them with both the slow equations and the fast equations coupled together, when the two FHN have different parameter values. To the best of our knowledge this is the first time it appears in symmetric coupling of identical equations. The mixed-mode oscillations arise close to the point where the two fold lines cross transversely. The general analysis of the dynamics in the neighbourhood of such a point is one of the task we intend to pursue in the near future.
@@
Because our results are based on normal hyperbolicity, they will persist under small symmetry breaking perturbations, as pointed out in Corollaries 4.4 and 4.5. However, asymmetric coupling, where one FHN is forcing the other, as in
@@
is not part of this scenario. Numerical evaluation of the Lyapunov spectrum of (17), presented in [14] for unidirectional coupling of two FHN through the fast equation, finds regions in the parameter space with frequency-locked solutions and other regions with chaotic behaviour. In Figure 10 we show a numerical solution of (17) that looks like a chaotic mixed-mode oscillation. This is a research direction we intend to pursue, but it is beyond the scope of the present article.
@@
Figure 10. Chaotic mixed-mode oscillations in (17), parameters $b=0$, $c=1.1505466726$, $k=0.820125$ and $\varepsilon=0.5$, initial condition $\left(x_{1},x_{2},y_{1},y_{2}\right)=\left(2,2,0,0\right)$. (A) - Time course for $x_{1}(t)$. (B) - Projection of the trajectory (blue) on the $\left(x_{1},y_{1}\right)$ plane, initial condition on the blue dot, red critical manifold $C_{0}$. (C) - Time course for $y_{1}(t)$.
(D) - Projection of the trajectory (blue) on the $\left(x_{1},x_{2}\right)$ plane, initial condition on the blue dot. (E) - Projection of the trajectory (blue) on the $\left(x_{2},y_{2}\right)$ plane, initial condition on the blue dot, red critical manifold $C_{0}$, equilibrium on the white dot. (F) - Time course for $x_{2}(t)$ (blue) and $y_{2}(t)$ (orange).
@@ Acknowledgments
The first and second authors had financial support from CMUP, member of LASI, (UIDP/00144/2020), financed by Fundação para a Ciência e a Tecnologia, Portugal (FCT/MCTES) through national funds. The third author has been supported by the Project CEMAPRE/REM (UIDB/05069/2020) financed by FCT/MCTES through national funds.

### v3

@@ 5.3. Double fold points
For (b) we have already established in Subsection 5.2 that in this case $\det DH\left(x_{1}^{*},x_{2}^{*}\right)>0$. From the expression (12) and assertion (a) it follows that when $c$ tends to $2\sigma/\sqrt{3}$ then $\det DH\left(x_{1}^{*},x_{2}^{*}\right)$ tends to 0. Therefore, for $c$ close to $2\sigma/\sqrt{3}$ we have $0<\det DH\left(x_{1}^{*},x_{2}^{*}\right)<1/4$ and since $\operatorname{tr}DH\left(x_{1}^{*},x_{2}^{*}\right)=1$, the point $\left(x_{1}^{*},x_{2}^{*}\right)$ is a node. ∎
@@ 6. Discussion
This article is an analysis of the dynamics of two identical FitzHugh–Nagumo equations symmetrically coupled through the slow equations. In this section we compare our results to findings by other authors using different types of coupling. Most authors only deal with the fast-to-fast coupling, probably motivated by its neurophysiological interpretation as an electrical synapse. The coupling described in the present article does not have a natural interpretation in neurophysiology but can be realised experimentally in electrical circuits, as is the case of the simultaneous slow-to-slow and fast-to-fast connections in Kawato et al. [21]. We have focused on synchrony, bistability and on mixed-mode oscillations generated by canards.
6.1. Synchrony and bistability
lixo
When two identical equations are coupled symmetrically there are always synchronous solutions, as is immediate from the symmetry. The interesting questions are whether the synchrony space contains non trivial dynamics, whether the synchronous solutions are asymptotically stable and hence visible, and whether approximate synchrony persists when the symmetry is broken. Similarly, coupling equations that have some additional symmetry always yields antisynchronous solutions to which analogous comments are applicable. Section 4 of this article is concerned with providing an answer to these questions in the present context.
We use the symmetry of the coupled equations to obtain in Theorem 4.2 attracting persistent synchronous solutions for all values of the parameters $b$ and $c$ in FHN and for sufficiently large coupling strengh $k$. In particular we show in Corollary 4.5 that if $b>1/4$ there are attracting synchronous periodic solutions even with negative coupling strength. We also show in Corollary 4.10 that approximate synchrony will persist if the symmetry is broken either by taking slightly different coupling constants or by a small change in the equations governing one of the cells.
@@ 6.1. Synchrony and bistability
Synchronised periodic solutions were also observed analytically by Campbell and Waite [3] and numerically in passing by Hoff et al. [20] when two FHN were coupled symmetrically and bidirectionally through the fast equations, but the latter do not report on persistence under symmetry breaking, their focus being on chaotic solutions. Kawato et al. [21] obtain both the synchrony and its persistence under symmetry breaking perturbations for two FHN with simultaneous slow-to-slow and fast-to-fast connection. Pedersen et al. [29] provide bifurcation diagrams in the synchrony plane for two FHN coupled symmetrically and bidirectionally through the fast equations, showing steady-state and Hopf bifurcations that create both stable and unstable limit cycles. Experimental results on synchrony of a radio engineering circuit that simulates two FHN with different values of the parameter $c$ coupled through the fast equations with time delay are presented in both Ponomarenko et al. [31] and Kulminskiy et al. [26], with the latter also containing numerical simulations. Both papers observe synchrony both in the symmetric case and in a large region in parameter space where the symmetry is broken. However, since their connection has a time delay, most of their findings are more similar to the antisynchrony situation discussed below.
Perfect and approximate synchrony are obtained analytically by Plotnikov and Fradkov [30] for two FHN differing on the parameter $c$ and coupled through the fast equations. They use a Lyapunov function to estimate the precision of synchronisation (cf. Definition 4.8) for large differences in the two values of the parameter $c$. For two identical FHN they also obtain perfect synchrony.
It is clear from all the results discussed above and ours that asymptotically stable synchrony arises from all types of linear coupling, even very far from the symmetric situation. An interesting future task is to extend the analysis done by Plotnikov and Fradkov [30] to a complete bifurcation study of all the types of coupling and of symmetry breaking.
@@
When the parameter $c$ is zero the individual FitzHugh–Nagumo equations becomes symmetrical under change of sign of both variables. This provides an additional symmetry to the coupled equations, with a flow-invariant fixed point subspace. From this we obtain Theorem 4.6 that implies the coexistence of different stable solutions. We show that the two sets of stable solutions persist when the additional symmetry is broken for small $c\neq 0$. The additional symmetry does not hold when the equations are coupled through the fast equations, so the bistability stated in Corollary 4.7 is a characteristic feature of the type of coupling analysed here.
Antisynchrony, in the sense of oscillations with a half period difference, was observed by Ponomarenko et al. [31] and by Kulminskiy et al. [26] in experiments with two FHN electrical circuits coupled through the fast equations with a time delay. Although their system has $c\neq 0$ and $b=0$, apparently the time delay induces a similar effect to the symmetry. They use a delay that seems to be close to half the period of the relaxation oscillations and the effect is observed both in two identical FHN and in the case of systems with different values of the parameter $c$ persisting for a large difference in parameter values.
@@
Bistability has been found in the work by Kawato et al. [21] coupling together the slow variables and the fast variables to each other in two FHN systems. When only the slow variables are coupled, they have obtained the coexistence of synchrony and antisynchrony through Hopf bifurcation in a parameter $I$ added to the fast equations. They also indicate that the bistability persists when the fast variables are also coupled with a small coupling constant.
Multistability has been found not associated to antisynchrony by Campbell and Waite [3] who considered electrical coupling of two FHN models through the fast equation. They show that when the magnitude of the coupling is strengthened, periodic orbits undergo several bifurcations (namely resonant Hopf-Hopf interactions) leading to the coexistence of a chaotic attractor with an attracting periodic orbit. This indicates that there might be other interesting dynamical structures that do not arise from the symmetry, a direction for further research. Note that neither [3] nor [21] used the fast-slow structure.
6.2. Mixed-mode oscillations
lixo
The symmetry in the model we have analysed limits the possible dynamic outcomes, and yet it still allows a number of interesting features. The presence of a canard induces small amplitude symmetry breaking oscillations before a solution converges to a synchronised periodic solution of large amplitude, shown in Figure 8. Oscillations that alternate between different amplitudes, called mixed-mode, have been associated to canards ever since they were first described in the van der Pol equations in Benoît et al. [1]. A detailed description of the way they appear in fast-slow systems can be found in Desroches et al. [7] and for the forced van der Pol equations in a recent article by J. Guckenheimer [17].
@@ 6.2. Mixed-mode oscillations
Santana et al. [33] report “transient chaos”, a complicated and long transient, in their numerical description of two FHN with different parameter values coupled through the fast equations. Canard-induced small amplitude transients were also found by Krisitiansen and Pedersen [22] in two identical FHN coupled through the fast equations. In their case the attracting periodic solution does not lie in the synchrony plane. The small amplitude transients arise from a canard at a cusp point in the critical manifold. In our case this manifold does not have cusps, so the origin of the transient oscillations is not the same, here they arise at a folded node – Proposition 5.6.
Sustained mixed-mode oscillations arising from canards, like those in Figures 9 and 10, are ubiquitous in coupled FHN. They appear in asymmetric coupling of the fast equation to the slow one in Doss-Bachelet et al. [10] and in Krupa et al. [24] where three different time scales are considered. Desroches et al. [6] find them in self-coupled FHN analogous to unidirectional coupling through the fast equation. Krupa et al. [23] find them with both the slow equations and the fast equations coupled together, when the two FHN have different parameter values. It is worth to read Yu et al. [37, pp. 364], where the authors investigate a Hopf bifurcation and the emergence of two types of mixed mode oscillations in a periodically-perturbed Bonhoeffer-van der Pol circuit. We also refer the reader to [5] where the fast-slow dynamics of a parametrically driven oscillator with triple-well potential has been studied. For some parameter values, the perturbed model combined with multiple frequency components may produce fast large oscillations and slow small oscillations clustered into the singular periodic attractor. Authors have shown that the critical manifold is folded with attracting and repelling branches separated by a fold. Such folded singularities can lead to oscillation behaviours with separated amplitudes – pairs of small-amplitude oscillations about stable manifold alternate with pairs of large-amplitude oscillations that occur due to jumps from one stable branch to the other.
By combining the dynamics within the synchrony and antisynchrony subspaces, we prove bistability in (2), where two types of solution coexist as hyperbolic attractors. They persist under small perturbation of the parameters. The symmetric coupling allows a manageable analysis of the system. Symmetry also plays the role of an “organizing center” – some hyperbolic structures in the full equivariant case persist in the asymmetric flow.
To the best of our knowledge this is the first time sustained mixed-mode oscillations appear in symmetric coupling of identical equations. It is possible that this is a consequence of the bistability arising from the additional symmetry of the equations: the large relaxation oscillations make excursions close to the antisynchrony plane, while the small amplitude canard oscillations remain in a neighbourhood of the synchrony plane. If this is the case, then it will not be found when two identical FHN are coupled symmetrically through the fast equations. The sustained mixed-mode oscillations arise close to the point where the two fold lines cross transversely as we established in Theorem 5.8. The general analysis of the dynamics in the neighbourhood of such a point is one of the tasks we intend to pursue in the near future.
6.3. One directional coupling
lixo
@@ 6.3. One directional coupling
Because our results are based on normal hyperbolicity, they will persist under small symmetry breaking perturbations, as pointed out in Corollaries 4.7 and 4.10. However, asymmetric coupling, where one FHN is forcing the other, as in
@@
is not part of this scenario. Numerical evaluation of the Lyapunov spectrum of (17), presented in [20] for unidirectional coupling of two FHN through the fast equation, finds regions in the parameter space with frequency-locked solutions and other regions with chaotic behaviour. In Figure 12 we show a numerical solution of (17) that looks like a chaotic mixed-mode oscillation. This is another research direction we intend to pursue, but it is beyond the scope of the present article.
@@
Figure 12. Chaotic mixed-mode oscillations in (17), parameters $b=0$, $c=1.1505466726$, $k=0.820125$ and $\varepsilon=0.5$, initial condition $\left(x_{1},x_{2},y_{1},y_{2}\right)=\left(2,2,0,0\right)$. These numerics illustrate the combination of large and small oscillations in a persistent and irregular way, possibly “chaotic” in a sense that deserves to be explored. (A) - Time course for $x_{1}(t)$. (B) - Projection of the trajectory (blue) on the $\left(x_{1},y_{1}\right)$ plane, initial condition on the blue dot, red critical manifold $C_{0}$. (C) - Time course for $y_{1}(t)$. (D) - Projection of the trajectory (blue) on the $\left(x_{1},x_{2}\right)$ plane, initial condition on the blue dot. (E) - Projection of the trajectory (blue) on the $\left(x_{2},y_{2}\right)$ plane, initial condition on the blue dot, red critical manifold $C_{0}$, equilibrium on the white dot. (F) - Time course for $x_{2}(t)$ (blue) and $y_{2}(t)$ (orange).
@@ Acknowledgments
The first and second authors had financial support from CMUP, member of LASI, (UIDP/00144/2020), financed by Fundação para a Ciência e a Tecnologia, Portugal (FCT/MCTES) through national funds. The third author has been supported by the Project CEMAPRE/REM (UIDB/05069/2020) financed by FCT/MCTES through national funds. The authors are also indebted to two reviewers for the suggestions which helped to improve the readability of this manuscript.